OpenStudy (anonymous):

Openstudy, A stoppered density bottle (pycnometer) has a volume of 5.000 ml. The empty bottle weighed 24.642 g. After some water was added, the mass was 28.297 g. Osmium powder was then added until the water level reached the top; the total mass was found to be 58.527 g. Neglecting temperature effects, and assuming the density of water to be 1.000 g/ml, find the density of osmium. This is the densest known substance at ordinary pressure.

OpenStudy (aaronq):

find the mass of the water (then find its volume using the density given) find the mass of the Os, then find it's volume (\(V_{Os}=5000 mL - V_{water}\)) then divide the mass of Os by it's volume. \(\rho=\dfrac{m}{V}\)

OpenStudy (anonymous):

how would i find the mass and volume of the water?

OpenStudy (aaronq):

Subtract the mass of the bottle from the combined mass of water and the bottle. when you find the mass you can use the density of water to find it's volume, use this equation: \(\rho=\dfrac{m}{V}\)

OpenStudy (anonymous):

alright, i got the volume by doing.. v= m/d which is 3.655 g/ml

OpenStudy (anonymous):

after calculating everything, i got the density of Osmium to be 22.475 g/ml. Is that correct? I wasn't sure about the units.

OpenStudy (goformit100):

First of all @gsethi2390 : A Warm Welcome to 'Open Study'. Please Read CoC (compulsory to read by all "Open Study" users) : http://openstudy.com/code-of-conduct